OpenStudy (anonymous):

A skateboarding ramp is 12in. high and rises at an angle of 17 degrees. How long is the base of the ramp? Round to the nearest inch.

OpenStudy (anonymous):

Starting with a good picture is crucial:  Now, since we know that the angles of any triangle add up to 180, we can add the other angle:  Now, we need to do a little trig. We could use the law of sines, which says: \[A/\sin a = B/\sin b\] Plugging in our values, we get: \[\frac{ 12 }{ \sin 17 }=\frac{ x }{ \sin 73 }\] Rearranging, we find: \[x= (\sin 73)(12) / \sin 17 = 39.25 inches\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem, glad I could help.